What is the present value of each of these three investments if the appropriate discount rate is 11 percent?


(Present value of an uneven stream of payments) You are given three investment alternatives to analyze. The cash flows from these three investments are as follows:

Investment
End of YearABC
1$1,000$2,000$5,000
22,0002,0005,000
33,0002,000(5,000)
4(4,000)2,000(5,000)
54,0005,00015,000

 What is the present value of each of these three investments if the appropriate discount rate is 11 percent?
Answer #1

Present value=Cash flows*Present value of discounting factor(rate%,time period)

A:

Present value=1000/1.11+2000/1.11^2+3000/1.11^3-4000/1.11^4+4000/1.11^5

=$4456.60(Approx)

B:

Present value=2000/1.11+2000/1.11^2+2000/1.11^3+2000/1.11^4+5000/1.11^5

=$9172.15(Approx)

C:

Present value=5000/1.11+5000/1.11^2-5000/1.11^3-5000/1.11^4+15000/1.11^5

=$10514.77(Approx).